OpenStudy (anonymous):

Please Help!!! write the equation in standard form y=3(x-4)^2-8

OpenStudy (anonymous):

1+1= you suck

OpenStudy (gorv):

standard y=ax^2 after shifting to some point(x1,y1) (y-y1)=a(x-x1)^2 will be standard form so move 8 on y side y+8=3(x-4)^2

OpenStudy (radar):

\[3(x ^{2}-8x+16)-8\]\[3x ^{2}-24x+48-8\]\[3x ^{2}-24x+40\]\[y=3x ^{2}-24x+40\]